# PLAICraft: Large-Scale Time-Aligned Vision-Speech-Action Dataset for Embodied AI

### v1

@@ Abstract
Advances in deep generative modelling have made it increasingly plausible to train human-level embodied agents. Yet progress has been limited by the absence of large-scale, real-time, multi-modal, and socially interactive datasets that reflect the sensory-motor complexity of natural environments. To address this, we present PLAICraft, a novel data collection platform and dataset capturing multiplayer Minecraft interactions across five time-aligned modalities: video, game output audio, microphone input audio, mouse, and keyboard actions. Each modality is logged with millisecond time precision, enabling the study of synchronous, embodied behaviour in a rich, open-ended world. The dataset comprises over 10,000 hours of gameplay from more than 10,000 global participants.111We have done a privacy review for the public release of an initial 200-hour subset of the dataset, with plans to release most of the dataset over time. Alongside the dataset, we provide an evaluation suite for benchmarking model capabilities in object recognition, spatial awareness, language grounding, and long-term memory. PLAICraft opens a path toward training and evaluating agents that act fluently and purposefully in real time, paving the way for truly embodied artificial intelligence.
@@ 1 Introduction
Embodied artificial intelligence (EAI) aims to build agents that perceive, act, and learn in dynamic environments through real-time sensory-motor interaction. As Brooks (1991) emphasized, such agents must “cope appropriately and in a timely fashion with changes in [their] dynamic environment” and “do something in [their] world; [they] should have some purpose in being.” These principles remain vital as research shifts toward agentic AI systems like Devin (AI, 2024) and Claude Code (Anthropic, 2024) that operate autonomously across extended tasks.
Yet most modern agents are disembodied in the sense that they interact with their environment indirectly, often through high-level text interfaces and asynchronous API calls. This limits their ability to engage in temporally grounded, perceptual-motor loops that characterize embodied intelligence. While simulated environments such as Habitat (Savva and et al., 2019), RoboTHOR (Deitke and et al., 2020), and ThreeDWorld (Gan and et al., 2021) support real-time, sensorimotor learning, they often lack socially interactive or linguistically rich settings. Within Minecraft, datasets like MineRL (Guss et al., 2019), OpenAI’s VPT (Baker and et al., 2022), and CraftAssist (Das and et al., 2019) have advanced perception-action and language learning. Most recently, MineDojo (Fan et al., 2022) introduced large-scale web-mined video-language data to support open-ended instruction following and knowledge transfer. However, these settings remain offline, non-interactive, heterogeneous, or lack real-time, social grounding—that is, the process of linking symbols to perceptual input, physical action, or shared environmental context.
We introduce PLAICraft, a dataset and platform designed for the study of real-time, socially situated, speech-interactive embodied agents. Built atop a modded multiplayer Minecraft server with a proximity-based voice chat plugin, PLAICraft records time-aligned streams of screen video, game output audio, player microphone input audio, keyboard, and mouse input with millisecond precision. Agents operate in persistent, open-ended worlds with other players where perception, speech, and action are tightly coupled in time. In its emphasis on rich, temporally grounded sensory-motor interaction, PLAICraft aligns with recent efforts to build datasets that support embodied learning in complex, naturalistic domains—such as autonomous driving benchmarks designed to capture diverse, real-world scenarios (Waymo, 2025; Zürn et al., 2024).
Figure 1 illustrates our approach: participants interact naturally while all modalities are recorded. Unlike most prior datasets, PLAICraft emphasizes dialogue, social context, and temporal embodiment, supporting the study of skills like object permanence, reactive speech, memory, and social reasoning. Agents must act with purpose amid uncertainty, social influence, and ambiguity.
@@
We have collected over 10,000 hours of data (Section 2), and publicly release a privacy-reviewed 200-hour subset (Section A.4) at https://blog.plaicraft.ai/2025/05/11/download-user-data/. We plan to release further data incrementally. The data collection platform is described in Section 3. Section 4 covers an evaluation suite designed to probe reasoning, memory, and communicative competence. We conclude in Section 5.
@@ 2 Plaicraft Dataset
We have chosen Minecraft as our simulated environment, like many previous datasets (Shah et al., 2021; Fan et al., 2022; Baker and et al., 2022; Das and et al., 2019), as its environment is sufficiently rich such that lessons learned within it can be transferred to the real world. In vanilla Minecraft, players navigate procedurally generated 3D worlds abundant with diverse terrains, resources to gather, tools to create, structures to construct, and novel discoveries to encounter. The world players interact with is also boundary-free, allowing unrestricted exploration. Minecraft provides neither explicit goals nor predefined storylines, rendering it uniquely suitable as a foundation for open-ended embodied AI research. Unlike many prior Minecraft datasets, which are mostly restricted to single-player mode with limited sets of tasks (Shah et al., 2021), limited modalities or nonsystematic internet-fetched videos (Fan et al., 2022), our PLAICraft dataset is unique in the following ways.
@@
Unbounded Multiplayer Social Environment: The open-ended nature of Minecraft inherently offers an unlimited spectrum of possible tasks. Our dataset further enhances this by employing a multiplayer environment where participants can socially interact with each other, and a proximity-based voice chat plugin for natural voice communication. This setup substantially enriches the complexity of achievable tasks, particularly enabling realistic social interactions. Figure 2 shows a wide range of solo and collaborative tasks players have performed on the server222For video examples, please check our blog website: https://blog.plaicraft.ai/..
@@
Our participants comprise Minecraft enthusiasts globally, spanning diverse age groups, genders, and experience levels. Self-reported demographic statistics of participants are illustrated in Figure 3. During each gameplay session, participants receive no specific task assignments or restrictions regarding interaction areas. Instead, they are encouraged to play as naturally as they would, thinking of this not as a research project but just regular Minecraft, subject only to basic behavioural guidelines promoting mutual respect. The virtual microphone on the VM is forced to be enabled, and we encourage our players to continuously talk, especially when they play interactively with others. In each gameplay session, players have a maximum of 4 hours of play time, thus, they can perform a wide range of different tasks and social interactions. For example, as shown in the left figure of Figure 4, the player engaged in a wide range of interactive activities from minute 1 to minute 22, but no interactive audio was recorded afterwards.
@@
Consequently, our dataset exhibits extensive dynamic variability. Solo player activities span a wide spectrum, beginning with basic actions such as moving with WASD keys or breaking a dirt block and progressing to moderately complex behaviours like mining, crafting, and combat. They extend further to highly advanced tasks, including defeating the Ender Dragon, constructing intricate redstone mechanisms, or even exploiting server bugs for personal benefit. Additionally, the dataset captures numerous social interactions, ranging from simple greetings and self-introductions to collaborative tasks involving team formation and cooperative gameplay. Notably, participants often form teams spontaneously, engaging in continuous interaction through collective base construction and various other tasks. Players between different teams interact with each other more cautiously and less frequently. Despite rigorous moderation efforts, occasional negative interactions such as conflicts, theft, and intentional property damage are present. Such behaviours are explicitly identified, and strong measures are implemented to prevent malicious participation.
@@
In addition to raw multimodal data, we provide encoded data representations generated by specialized autoencoders (detailed in Section 3). The right figure of Figure 4 illustrates the temporal alignment of these encoded formats. Given the differing sampling rates of video encodings (10 frames per second) and audio encodings (75 frames per second), each video frame corresponds to 7.5 audio frames. To resolve this fractional correspondence, we define the minimal encoding unit as two consecutive video frames (200 milliseconds), forming the basic processing unit used by our dataloader (see Section A.12).
@@ 3 Data Collection
Gathering gameplay data often involves developing specialized platforms tailored to specific games, each requiring unique strategies for user engagement and data collection. MineRL (Guss et al., 2019), for instance, utilizes a server and plugin system that players can access for gameplay. However, this platform mainly captures data within a predefined set of actions and tasks. On the other hand, the data set provided by MineDojo (Fan et al., 2022) is crawled from disjoint, pre-existing internet platforms such as YouTube, Wikipedia, and Reddit. While it contains a wide array of gameplay styles and strategies, it also introduces significant variability in data quality and distribution. The PLAICraft data collection platform addresses these limitations by standardizing the data collection process, ensuring the capturing of time-aligned video, game output audio, microphone input audio, keyboard and mouse action data while providing participants the freedom to naturally and collaboratively engage within a multi-player Minecraft setting.
@@ 3.1 Platform Setup
As illustrated in Figure 6, our platform is constructed mainly using Amazon Web Services (AWS). The central component is a remote-controlled virtual machine (VM) running as an AWS Ubuntu EC2 instance. Each session is hosted on a dedicated VM instantiated from an identical AMI image, ensuring secure, isolated, independent and uniform data recording, with automatic data upload upon session completion. The VM is equipped with Amazon NICE DCV, a high-performance remote desktop and application streaming protocol, in conjunction with an NVIDIA T4 GPU. In combination, they provide a seamless Minecraft gaming experience featuring shader-enhanced visuals and high-quality, consistent and smooth video recordings. The platform is designed to immerse participants in gameplay upon connection to the EC2 instance, with the session concluding when the player stops gameplay. We ensure this by disabling any UI interaction players can perform with the desktop other than the Minecraft window. The Minecraft client launches in full-screen mode and cannot be minimized. Upon instance initialization, auxiliary monitoring scripts track both player connectivity and window focus; if they detect that the player has disconnected or shifted focus away from the Minecraft window, the scripts automatically terminate the EC2 instance and upload the recorded data to an Amazon S3 bucket.
Leveraging the Amazon DCV web frontend alongside AWS CloudFront, participants can effortlessly engage with the platform from any browser on any computer with a Chromium-based browser by logging into plaicraft.ai. We deployed our infrastructure at various AWS data center locations. Upon login, an instance is dynamically provisioned at the nearest available data center, ensuring stable connection speeds and low latency. Consequently, players experience consistently high-performance gameplay at a stable 60 FPS irrespective of their local hardware capabilities. Our platform offers a similar experience to popular cloud gaming services such as GeForce NOW and Xbox Cloud Gaming. Moreover, as all recording processes occur within the secure environment of the virtual machine started from the same AMI image, participant privacy is fully protected, and the dimensionality and semantics of all modalities’ data collected are uniform and fixed. Nonetheless, NICE DCV’s constraints allow mismatched screen aspect ratios to appear as black borders in the recorded video, while the audio remains susceptible to background noise and the varying quality of participants’ microphones.
@@ 3.2 Minecraft Configurations
To ensure game stability and compatibility with various modifications, we fixed the Minecraft version to 1.19.4 and retained most vanilla features. Players exclusively engage in survival mode with normal difficulty settings, and all three standard worlds (the Overworld, Nether, and the End) are fully accessible. To encourage social interactions, teleportation capabilities and a global chat interface are provided.
We incorporated multiple custom-developed and third-party plugins to enhance players’ experience. Central to facilitating game play with over 10,000 participants is our AutoJoin plugin, which allows multiple participants to use a shared Minecraft license while preserving individual game states. Since providing unique licenses for all participants is impractical, this plugin enables us to maintain a minimal number of licenses while supporting an unlimited number of participants. The Autojoin plugin also automatically relocates first-time players to a safe position near the world spawn. To prevent localized overcrowding, we periodically shift the spawn point to new regions.
@@
Another critical enhancement is the Simple Voice Chat plugin, a Spigot-based modification enabling immersive, proximity-based 3D voice communication within the game. Players can communicate naturally using microphones, with audio volume dynamically adjusting based on in-game distance and relative positioning. This feature substantially enriches player interactions, elevating the dataset’s complexity and realism to a level comparable with real-world interaction datasets.
@@
Finally, the inclusion of the Complementary Shader Pack (Unbound) significantly improves visual detail and complexity closer to real-world environments, enabling features like real-time shading, realistic clouds, water flow with reflections and a number of other high-quality effects. This add-on helped us attract a lot more participants and makes the dataset’s visual space more realistic. A comprehensive list of applied plugins is provided in the Section A.6.
3.3 Recording Software
Video & Audio Video recordings are captured using Open Broadcaster Software (OBS) running on the EC2 instance. Videos are recorded at a constant frame rate of 30 FPS with a resolution of 1280x720, encoded using the H.264 codec. OBS is specifically configured to capture only the Minecraft application window in full-screen mode, ensuring that all recorded frames pertain solely to gameplay. Audio streams are simultaneously recorded by OBS, separately capturing user microphone input and game audio output. Each audio stream is recorded at a constant sampling rate of 48 kHz in stereo format, encoded with the AAC codec.
Mouse & Keyboard Mouse and keyboard inputs are precisely captured using a custom-developed Forge mod integrated directly into the client-side Minecraft application. This method provides superior temporal accuracy, capturing server input event processing time compared to traditional system-level logging that is unaware of the Minecraft server. Input events for both mouse and keyboard are recorded on an event-driven basis with millisecond precision, maintaining a maximum polling rate of 100 Hz, resulting in a minimum interval of 10 ms between consecutive data points. Additionally, comprehensive system-level logging using pyxhook is implemented as a secondary data collection measure. These system-level logs, although less precise in timestamp accuracy, serve primarily to validate the integrity and completeness of the mod-generated input data.
3.4 Data Filtering and Pre-processing
We implemented comprehensive filtering and pre-processing strategies to maintain the integrity and quality of our dataset. Initially, real-time filtering was deployed on the Minecraft server through dedicated plugins for profanity detection and grief prevention, mitigating malicious behaviours and inappropriate content during gameplay. Post-recording, further pre-processing steps were conducted to ensure that all collected data pertained strictly to gameplay activities. All modalities underwent precise trimming aligned to gameplay start times, achieving synchronization at millisecond precision. For spoken audio streams, noise filtering using Voice Activity Detection (VAD) was employed alongside speaker diarization, ensuring audio recordings contained clear, isolated speech segments from a single speaker at any given time. Additionally, we also provide the transcription with word-level timestamps for both audio streams by running WhisperX (Bain et al., 2023) on the denoised version of the two audio tracks.
To optimize data handling and reduce GPU resource demand during model training, we pre-encoded the collected modalities. Video data were encoded using the sdxl-vae-fp16-fix, a variant of the SDXL Variational Autoencoder (VAE) (Podell et al., 2023) optimized for 16-bit floating-point precision, selected for its superior ability to preserve textual details in the game’s graphical user interface. Audio data of speaking and hearing was individually encoded using the Encodec 24 kHz model (Défossez et al., 2022), a state-of-the-art causal neural audio codec model with residual vector quantizers pretrained on a variety of audio data. Additionally, keyboard and mouse click data were compressed through a custom-trained temporal convolutional autoencoder with GRU layers to achieve dimensionality reduction. Given the inherently low dimensionality of mouse movement data, we opted against encoding it, choosing instead to simply apply global normalization.
3.5 Promotion Effort & Data Collection History
Hiring expert gamers to provide large amounts of demonstration data is usually costly and infeasible (DeepMind, 2019; Fan et al., 2022). However, being the most popular game in the world in terms of sales (Wikipedia, 2025), Minecraft has a huge player community globally. All we need to do is reach out to this community and attract them to our platform. With an extensive media outreach and promotion efforts, see Section A.7, we have gradually formed a large community ourselves with Minecraft enthusiasts and experts. Figure 7 demonstrates our data collection history.
3.6 Ethical Considerations
This research ensures the safety and privacy of participants through the guidance of UBC’s Office of Research Ethics333https://researchethics.ubc.ca/. Study procedures are reviewed routinely by the Behavioural Research Ethics Board (BREB); and have been found acceptable on ethical grounds for research involving human subjects, as well as deemed minimal risk to participants. Participants’ data is collected with their consent, or that of their guardian, and any reasonably identifiable aspects of this data are anonymized systematically. We intend to release access to our full dataset with time, once participants have been given the opportunity to withdraw.

### v2

@@ Abstract
Advances in deep generative modeling have made it increasingly plausible to train human-level embodied agents. Yet progress has been limited by the absence of large-scale, real-time, multi-modal, and socially interactive datasets that reflect the sensory-motor complexity of natural environments. To address this, we present PLAICraft, a novel data collection platform and dataset capturing multiplayer Minecraft interactions across five time-aligned modalities: video, game output audio, microphone input audio, mouse, and keyboard actions. Each modality is logged with millisecond time precision, enabling the study of synchronous, embodied behaviour in a rich, open-ended world. The dataset comprises over 10,000 hours of gameplay from more than 10,000 global participants. Alongside the dataset, we provide an evaluation suite for benchmarking model capabilities in object recognition, spatial awareness, language grounding, and long-term memory. PLAICraft opens a path toward training and evaluating agents that act fluently and purposefully in real time, paving the way for truly embodied artificial intelligence.
@@ 1 Introduction
Embodied artificial intelligence (EAI) aims to build agents that perceive, act, and learn in dynamic environments through real-time sensory-motor interaction. As Brooks [6] emphasized, such agents must “cope appropriately and in a timely fashion with changes in [their] dynamic environment” and “do something in [their] world; [they] should have some purpose in being.” These principles remain vital as research shifts toward agentic AI systems like Devin [1] and Claude Code [2] that operate autonomously across extended tasks.
Yet most modern agents are disembodied in the sense that they interact with their environment indirectly, often through high-level text interfaces and asynchronous API calls. This limits their ability to engage in temporally grounded, perceptual-motor loops that characterize embodied intelligence. While simulated environments such as Habitat [23], RoboTHOR [11], and ThreeDWorld [15] support real-time, sensorimotor learning, they often lack socially interactive or linguistically rich scenarios. Minecraft datasets like MineRL [17], OpenAI’s VPT [5], and CraftAssist [8] have driven progress in perception–action coupling and language learning. Most recently, MineDojo [14] introduced large-scale web-mined video-language data to support open-ended instruction following and knowledge transfer. However, these settings remain offline, non-interactive, heterogeneous, or lack real-time, social grounding—that is, the process of linking symbols to perceptual input, physical action, or shared environmental context.
We introduce PLAICraft, a dataset and platform designed for the study of real-time, socially situated, speech-interactive embodied agents. Built atop a modded multiplayer Minecraft server with a proximity-based voice chat plugin, PLAICraft records time-aligned streams of screen video, game output audio, player microphone input audio, keyboard, and mouse input with millisecond precision. Agents operate in persistent, open-ended worlds with other players where perception, speech, and action are tightly coupled in time. In its emphasis on rich, temporally grounded sensory-motor interaction, PLAICraft aligns with recent efforts to build datasets that support embodied learning in complex, naturalistic domains—such as autonomous driving benchmarks designed to capture diverse, real-world scenarios [31, 35].
Fig.˜1 illustrates our approach: participants play in our virtual environment and interact with other players naturally while all modalities are recorded. Unlike most prior datasets, PLAICraft emphasizes dialogue, social context, and temporal embodiment, supporting the study of skills like object permanence, reactive speech, memory, and social reasoning. Agents must act with purpose amid uncertainty, social influence, and ambiguity.
@@
We have collected over 10,000 hours of data from over 10,000 players around the world. We introduce the PLAICraft dataset in Sec.˜2. The PLAICraft data collection platform is described in Sec.˜3. Sec.˜4 covers an evaluation suite designed to probe reasoning, memory, and communicative competence. We conclude in Sec.˜6 and Sec.˜8.
@@ 2 Plaicraft Dataset
We have chosen Minecraft as our environment, like many previous datasets [25, 14, 5, 8], because its rich, open-ended dynamics make learned behaviours transferable to real-world scenarios. In vanilla Minecraft, players navigate procedurally generated 3D worlds abundant with diverse terrains, resources to gather, tools to create, structures to construct, and novel discoveries to encounter. The world players interact with is also boundary-free, allowing unrestricted exploration. Minecraft provides neither explicit goals nor predefined storylines, rendering it uniquely suitable as a foundation for open-ended embodied AI research. Unlike many prior Minecraft datasets, which are mostly restricted to single-player mode with limited sets of tasks [25], limited modalities or non-systematic internet-fetched videos [14], our PLAICraft dataset is unique in the following ways.
@@
Unbounded Multiplayer Social Environment: The open-task nature of Minecraft inherently offers an unlimited spectrum of possible tasks. Our dataset further enhances this by employing a multiplayer environment where participants are encouraged to socially interact with each other, and a proximity-based voice chat plugin for natural voice communication. This setup substantially enriches the complexity of behaviours, particularly enabling realistic social interactions. Figure 2 shows a wide range of solo and collaborative activities players have performed on the server.
@@
Our participants comprise Minecraft enthusiasts globally, spanning diverse age groups, genders, and experience levels. Self-reported demographic statistics of participants are illustrated in Figure 3. During each gameplay session, participants receive no specific task assignments or restrictions regarding interaction areas. Instead, they are encouraged to play naturally, thinking of this not as a research project but just regular Minecraft, subject only to basic behavioural guidelines promoting mutual respect. The virtual microphone on the VM is forced to be enabled, and we encourage our players to continuously talk, especially when they play interactively with others. In each gameplay session, players have a maximum of 4 hours of play time; thus, they can perform a wide range of different activities. For example, as shown in Figure˜4, the player engaged in a wide range of interactive activities from minute 1 to minute 22, and from minute 22 to the end, the player only performed solo activities.
@@
Consequently, our dataset exhibits extensive dynamic variability. Solo player activities span a wide spectrum, beginning with basic actions such as moving with WASD keys or breaking a dirt block and progressing to moderately complex behaviours like mining, crafting, and combating. They extend further to events that require advanced skills, including defeating the Ender Dragon, constructing intricate redstone mechanisms, or even exploiting server bugs for personal benefit. Additionally, the dataset captures numerous social interactions, ranging from simple greetings and self-introductions to collaborative activities involving team formation and cooperative gameplay. Notably, participants often form teams spontaneously, engaging in continuous interaction through collective base construction and various other activities. Players between different teams interact with each other more cautiously and less frequently. Despite rigorous moderation efforts, occasional negative interactions such as conflicts, theft, and intentional property damage are present. Such behaviours are explicitly identified, and strong measures are implemented to prevent malicious participation.
@@
In addition to raw multimodal data, we provide encoded data representations generated by specialized autoencoders (detailed in Section 3). Figure˜5 illustrates the temporal alignment of these encoded formats. Given the differing sampling rates of video encodings (10 frames per second) and audio encodings (75 frames per second), each video frame corresponds to 7.5 audio frames. To resolve this fractional correspondence, we define the minimal encoding unit as two consecutive video frames (200 milliseconds).
@@ 3 Data Collection
Gathering gameplay data often involves developing specialized platforms tailored to specific games, each requiring unique strategies for user engagement and data collection. MineRL [17], for instance, utilizes a server and plugin system that players can access for gameplay. The data set provided by MineDojo [14] is crawled from pre-existing internet platforms such as YouTube, Wikipedia, and Reddit. The PLAICraft data collection platform standarizes the data collection process, ensuring the capturing of time-aligned video, game output audio, microphone input audio, keyboard and mouse action data while providing participants the freedom to naturally and collaboratively engage within a multi-player Minecraft setting.
@@ 3.1 Platform Setup
As illustrated in Figure 7, our platform is built primarily on Amazon Web Services (AWS). The core of the architecture is a set of AWS Lambda functions that serve as the backend. These Python functions handle interactions with various AWS resources; for example, some functions read from and write to Amazon DynamoDB, the database service where we store user information and gameplay status. Each Lambda function is fronted by Amazon API Gateway, which exposes a secure endpoint for external invocation.
Users interact exclusively with the front end, which is a simple website accessible in any Chromium-based browser. The front end is hosted in Amazon S3 and delivered to users via Amazon CloudFront, a global content delivery network. Players log in with their email address or phone number and are then redirected to an Amazon EC2 instance where they can play the game.
The EC2 instance is the virtual server on which we run the Minecraft game and associated data-collection software. We pre-package an Amazon Machine Image (AMI) from an instance with all software and environment configuration. We configured the instance to immerse participants in gameplay by disabling any UI interactions other than the Minecraft window. Upon user login, a dedicated g5.2xlarge instance will be spawned from the same AMI and bound exclusively to that user. This ensures secure, isolated and uniform data recording. The Minecraft client launches in full-screen mode and cannot be minimized. We also run auxiliary monitoring scripts that track the player’s connectivity and window focus; if they detect that the player has disconnected or shifted focus away from the Minecraft window, the scripts automatically terminate the EC2 instance and upload the recorded data to S3.
The instance is equipped with NVIDIA A10G Tensor Core GPUs so that our GPU-demanding game and software can run smoothly. The instance’s display is streamed to the user through NICE DCV, a high-performance remote display protocol, at up to 60 FPS and up to 2K resolution. Our platform offers a similar experience to popular cloud gaming services such as GeForce NOW and XBOX Cloud Gaming. Moreover, since recording occur within the secure environment of the virtual machine, participant privacy is fully protected.
@@ 3.2 Minecraft Configurations
We retained most features of vanilla Java Minecraft. Players exclusively engage in survival mode with normal difficulty settings, and all three standard worlds (the Overworld, Nether, and the End) are fully accessible. To encourage social interactions, teleportation capabilities and a global chat interface are provided.
We used a combination of custom and third-party plugins to enhance the player experience. In particular, because buying unique game licenses to every participant would be impractical, we developed the Autojoin plugin, which saves and restores each player’s state (inventory, position, level, etc.) in our DynamoDB backend rather than binding it to a specific license. This decoupling allows us to reuse a limited pool of licenses across non-concurrent players, enabling the platform to scale to more than 10,000 participants.
@@
The most critical enhancement is the integration of the Simple Voice Chat plugin, a Spigot-based modification enabling immersive, proximity-based 3D voice communication within the game. Players can communicate naturally using microphones, with audio volume dynamically adjusting based on in-game distance and relative positioning. This feature substantially enriches player interactions, elevating the dataset’s complexity and realism to a level comparable with real-world interaction datasets.
@@
Finally, the inclusion of the Complementary Shader Pack (Unbound) significantly improves visual detail and complexity closer to real-world environments. A comprehensive list of applied plugins is provided in the Section˜A.3.
3.3 Increasing Social Interactions
Social interaction is a central focus of our dataset, and we employ multiple mechanisms to encourage social engagement. We enable a persistent multiplayer environment, integrate a voice-chat plugin, and deploy curated mods that lower friction for collaboration and communication. We further provide UI/UX features in both the web client and in-game interfaces that streamline inviting friends, forming parties, and coordinating play.
We also run live transcription using WhisperX to process players’ speech audio and their proximate hearing audio during gameplay. The resulting transcripts are analyzed by a large language model to detect and characterize meaningful conversations with other players. Based on the quality and quantity of such interactions, we allocate additional solo play time as an incentive, thereby reinforcing sustained, socially rich behavior.
3.4 Data Recording & Pre-processing
We synchronously capture all modalities on the EC2 instance using dedicated software components, as shown in Table˜1. We use the exact timestamp at which OBS’s ffmpeg muxer begins writing the video file as the global start time and align all modalities to this reference. We then apply filtering to remove noisy data and harmful content, after which the data are pre-encoded using specialized autoencoders to reduce computational burden during model training. The detailed pipeline is described in Section˜A.6 and Section˜A.7.
| Modality | Software | Raw format | Encoding method | Encoded format |
| Video | OBS | 30 FPS, 1280720, H.264 | sdxl-vae-fp16-fix [22] | 10FPS, (C=4, W=160, H=96) |
| Speaking audio | OBS | 48 kHz, stereo, AAC | Encodec 24 kHz [défossez2022highfidelityneuralaudio] | 75Hz, (D=128) |
| Hearing audio | OBS | 48 kHz, stereo, AAC | Encodec 24 kHz [défossez2022highfidelityneuralaudio] | 75Hz, (D=128) |
| Mouse movement | Forge mod | 100 Hz, (D=2) | Normalization | 100 Hz, (D=2) |
| Key presses | Forge mod | 100 Hz, (D=79) | Custom autoencoder | 50Hz, (D=16) |
3.5 Data Collection History
We do not hire experts to play since Minecraft already has the largest player number globally. With extensive media outreach and promotion efforts, we have formed a large community of voluntary players consists of Minecraft enthusiasts and experts. Figure 8 demonstrates our data collection and player base development history.